OpenStudy (anonymous):

For a sinusoidal function in the form y = A sin Bx + C or y = A cos Bx + C, the period is 1, the minimum is negative 7, the equation of the midline is y = minus 2, the y-intercept is 3, and A > 0. What is the equation of the function? y = 5 sin (2pi x) minus 2 y = 9 sin (2pi x) minus 2 y = 5 cos (2pi x) minus 2 y = 9 cos (2pi x) minus 2

OpenStudy (anonymous):

I AM SEEING DIAMONDS???

OpenStudy (anonymous):

�������������

OpenStudy (anonymous):

idk why that literally just happened when i added you

OpenStudy (anonymous):

still seeing diamonds lol

OpenStudy (anonymous):

I can see till period is 1

OpenStudy (anonymous):

the minus sign is making the diamonds appear

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

they all have period 1 since \(\frac{2\pi}{2\pi}=1\)

OpenStudy (anonymous):

misty can you see the question??

OpenStudy (misty1212):

the minimum of sine and cosine is -1, so choose one that has a 5 at the front and a \(-2\) at the end

OpenStudy (anonymous):

now we're talking

OpenStudy (misty1212):

we have narrowed it down to A and C

OpenStudy (anonymous):

y intercept is three

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

substitue x=0 and check in which y=3

OpenStudy (anonymous):

the ans that gives you y=3 is the one

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

easy peesy Japaneesy

OpenStudy (anonymous):

can you now tell whats the ans??...A/B/C/D?? try

OpenStudy (misty1212):

it is probably C because it is always C

OpenStudy (anonymous):

so its C o.o

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

cheerio.:)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

btw ...I really need tough ones!!.this was good tho!

OpenStudy (anonymous):

lol ill try

OpenStudy (anonymous):

k then A'ight